OpenStudy (peterg2001):

please help A gift basket has 2 pounds of cookies, 3/5 pound of fudge, and 1 1/4pounds of nuts. What is the total weight of the items in the basket?

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

2+3/5+11/4=5 7/20

OpenStudy (peterg2001):

the answer is 3 17/20?

OpenStudy (anonymous):

yes it is 3(17/20)

OpenStudy (alexandervonhumboldt2):

no 5 7/20

OpenStudy (peterg2001):

wrong

OpenStudy (anonymous):

If you don't believe me type it into your calculator and double check

OpenStudy (anonymous):

but just for you I will work it out

OpenStudy (anonymous):

so first we find the least common denominator between 5 and 4 which in this case is 20

OpenStudy (alexandervonhumboldt2):

2+12/20+55/20=(40+12+55)/20=107/20=5 7/20

OpenStudy (anonymous):

and so we multiply so that each fraction is divided by 20

OpenStudy (peterg2001):

could be but i dont have that 14 3/4 -14 3/4 88 3/4 88 11/12

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

i dont know i have a calculator and it says 5 7/20

OpenStudy (anonymous):

2*20= 40 so we have 40/20 4*5=20 so we multiply top and bottom by 4 and we get 12/20 5*4=20 so we multiply top and bottom by 5 and we have 1(5/20)

OpenStudy (alexandervonhumboldt2):

why *?

OpenStudy (alexandervonhumboldt2):

add

OpenStudy (peterg2001):

so the answer is 14 3/4?

OpenStudy (alexandervonhumboldt2):

i mdont know

OpenStudy (anonymous):

now 1=20/20 so we add that and we have\[\frac{ 40 }{ 20 }+\frac{ 12 }{ 20 }+\frac{ 25 }{ 20 }\]

OpenStudy (alexandervonhumboldt2):

www.mathway.com try this

OpenStudy (anonymous):

this equals\[\frac{ 77 }{ 20 }\]

OpenStudy (anonymous):

20 goes into 77, 3 times, and we have a remainder of 17

OpenStudy (alexandervonhumboldt2):

your last fraction is wrong

OpenStudy (alexandervonhumboldt2):

11*5=55 not 25!!

OpenStudy (anonymous):

it's not 11 in the original problem. It's 1

OpenStudy (anonymous):

it's\[1\frac{ 1 }{ 4 }\]

OpenStudy (alexandervonhumboldt2):

oh well then let me count

OpenStudy (alexandervonhumboldt2):

3 17/20